OpenStudy (anonymous):

A triangle drrawn with angles of 40 degrees at the top and 65 degrees on the far right with a measurement of 72 on one side(right side line) sits next too a triangle of a 40 degree angle at the top with a line measurement of 52 on the left side and 45 line measurement on the right side. What would the length of all the missing sides and missing angles be?

OpenStudy (anonymous):

In the first triangle the last angle is given by 180-40-65=...

OpenStudy (amistre64):

you can use a mixture of the law of cosine and the law of sines to determine these...

OpenStudy (amistre64):

2nd triangle: a=52; C=40 b=45 Find B (bottom left angle), A (bottom right angle), and c (bottom line) c^2 = a^2 + b^2 -2ab cosC c^2 = 52^2 + 45^2 -2(52)(45)(cos(40)) c^2 =

OpenStudy (amistre64):

c^2 = 4729 - 4680(cos(40))

OpenStudy (amistre64):

c^2 = 1143.9120062031827952528023954006 c = 33.8218 if i did it right....

OpenStudy (amistre64):

sinA sinB sinC ---- = ---- = ---- a b c

OpenStudy (anonymous):

Thank you very much